(1+tan1)(1+tan2)(1+tan45)

Grade:12th Pass

1 Answers

Gurwinder Kaur
12 years ago

= (1+ tan2+tan1+tan1tan2)(1+tan45)      ;As tan45 = 1

= (1+ tan2+tan1+tan1tan2) (1+1)

= 2 (1+ tan2+tan1+tan1tan2)

= (2+ 2tan2+2tan1+2tan1tan2)

plz approve if helpful...